# Odd dynamics of passive objects in a chiral active bath

### v1

@@ Abstract
When submerged in a chiral active bath, a passive object becomes a spinning ratchet imbued with odd transport properties. In the adiabatic limit of a massive object, we derive the most general Langevin dynamics for a rigid body in a chiral active bath, with odd diffusion and odd mobility connected by an Einstein relation, and numerically explore the breakdown of these predictions beyond the adiabatic limit. We show that the irreversibility of a massive object increases as its symmetry decreases: a disk exhibits an effective equilibrium dynamics, while a rod admits distinct translational and rotational temperatures, and a wedge is fully out of equilibrium. Conversely, this departure from equilibrium can be read in the universal far-field currents and density modulations of the bath, which we measure numerically and derive analytically.
@@
A microscopic object in a fluid bath inherits its dynamics from collisions with the bath particles. When the bath is in equilibrium, the object obeys the Einstein relation and exhibits Boltzmann statistics. Active baths, by contrast, allow a richer set of phenomena and have thus attracted a lot of attention Wu and Libchaber (2000); Loi et al. (2008); Underhill et al. (2008); Leptos et al. (2009); Dunkel et al. (2010); Kurtuldu et al. (2011); Foffano et al. (2012); Miño et al. (2013); Morozov and Marenduzzo (2014); Burkholder and Brady (2017); Chaki and Chakrabarti (2018); Kanazawa et al. (2020); Knežević and Stark (2020); Reichert et al. (2021); Maes (2020); Granek et al. (2022); Solon and Horowitz (2022); Sorkin et al. (2024). Notable among these is ratchet motion Magnasco (1993); Van den Broeck et al. (2004); Angelani et al. (2011); Gnoli et al. (2013); Reichhardt and Reichhardt (2017), which demonstrates how an active bath can power microscopic engines Krishnamurthy et al. (2016); Zakine et al. (2017); Pietzonka et al. (2019); Fodor and Cates (2021). When asymmetric gears are inserted into an active bacterial bath Di Leonardo et al. (2010); Sokolov et al. (2010); Anand et al. (2024), rotational ratchet motion arises due to broken time-reversal symmetry of the bath and broken parity symmetry (i.e. chirality) of the object.
@@
A chiral active bath, however, breaks both these symmetries by itself, so that even symmetric objects become rotational ratchets Li et al. (2023) imbued with odd diffusivity Hargus et al. (2021); Kalz et al. (2022) and odd mobility Reichhardt and Reichhardt (2019); Poggioli and Limmer (2023). While extensive analytical and numerical results are available for achiral baths, little is known theoretically of the dynamics of passive objects embedded in chiral active fluids. The latter are however abundant in both biological Petroff et al. (2015); Drescher et al. (2009); Tan et al. (2022) and synthetic active systems Bililign et al. (2022). This raises two key questions: how are the emerging properties of the object (and the bath) interrelated, and how are they influenced by the object shape?
@@
In this Letter, we answer these questions by constructing a general theory to describe the dynamics of a massive object embedded in a chiral active bath. This allows us to show that the object dynamics become increasingly rich as its symmetry decreases. Specifically, we consider the disk, rod, and wedge depicted in Fig. 1. For the disk, we find an effective equilibrium description for the positional degrees of freedom, accompanied by circulating currents in momentum space. The rod already departs thermal equilibrium by featuring unequal effective temperatures for its rotational and positional degrees of freedom. The wedge fully departs from any effective equilibrium and behaves as both a translational and rotational ratchet. Finally, we show how the object shape influences the structure and flow of the bath itself through the generation of density modulations and currents that increase as the object symmetry decreases. Numerical details are provided in the End Matter.
@@
$M\ddot{\bm{R}}=\dot{\bm{P}}=\text{\bf F}\,,\quad I\ddot{\Theta}=\dot{L}=\Gamma\,,$ (1)
@@
where F and $\Gamma$ are the total force and torque due to the bath particles. The latter are modeled as chiral active Brownian particles, whose positions $\bm{r}_{i}$ and orientations $\theta_{i}$ evolve as
$\gamma\dot{\text{\bf r}}_{i}=\text{\bf F}_{i}+f_{0}\text{\bf u}_{i}+\sqrt{2\gamma D_{t}}\bm{\eta}_{i}\,,\quad\dot{\theta}_{i}=\omega_{0}+\sqrt{2D_{r}}\xi_{i}\,.$ (2)
@@
Here, $\gamma$ is the substrate friction, $f_{0}$ is the active force oriented along the director $\text{\bf u}_{i}=[\cos(\theta_{i}),\sin(\theta_{i})]^{T}$, and object-bath interactions are reciprocal, $\sum_{i}\text{\bf F}_{i}=-\text{\bf F}$. The bath displays two active lengthscales: the persistence length $\ell_{p}=\frac{f_{0}}{D_{r}\gamma}$ and gyroradius $\ell_{g}=\frac{f_{0}}{|\omega_{0}|\gamma}$.
Figure 1: Passive objects inherit chiral dynamics from the bath. A disc (a) exhibits odd diffusivity $D_{\perp}$, a rod (b) additionally rotates with a ratchet velocity $\Omega$, and a passive wedge (c) additionally translates with ratchet velocities $v_{\parallel}$ and $v_{\perp}$.
@@
If the motion of the object is very slow relative to that of the bath, we can work in the adiabatic limit Van Kampen and Oppenheim (1986); Granek et al. (2020); Solon and Horowitz (2022); Jayaram and Speck (2023) where the probability distribution of the full system factorizes as $\rho({\bm{R}},{\bm{P}},\Theta,L,\text{\bf r}^{N},\text{\bf u}^{N},t)\approx\rho_{\text{o}}({\bm{R}},{\bm{P}},\Theta,L,t)\rho_{\text{b}}(\text{\bf r}^{N},\text{\bf u}^{N}|{\bm{R}},\Theta)$. This leaves all time dependence in the probability of the object, while the bath relaxes instantaneously, following adiabatically the object configuration. This separation of timescales is characterized by the dimensionless parameter $\epsilon=\sqrt{\gamma/D_{r}M}$, which is small in the limit of massive objects. Following established procedures Van Kampen and Oppenheim (1986); Solon and Horowitz (2022), we integrate out the bath degrees of freedom to obtain underdamped Langevin dynamics for the object:
@@
$\begin{split}\begin{bmatrix}\dot{\bm{P}}\\ \dot{L}\end{bmatrix}&=\begin{bmatrix}\langle\text{\bf F}\rangle_{\text{b}}\\ \langle\Gamma\rangle_{\text{b}}\end{bmatrix}-\underbrace{\begin{bmatrix}\bm{\zeta}_{{\bm{P}}{\bm{P}}}&\bm{\zeta}_{{\bm{P}}L}\\ \bm{\zeta}_{L{\bm{P}}}&\zeta_{LL}\end{bmatrix}}_{\bm{\zeta}}\begin{bmatrix}\frac{1}{M}{\bm{P}}\\ \frac{1}{I}L\end{bmatrix}+\underbrace{\begin{bmatrix}\bm{\xi}_{P}(t)\\ \xi_{L}(t)\end{bmatrix}}_{\bm{\xi}}\,.\end{split}$ (3)
@@
The bath produces three effects on the object: a mean force $\langle\text{\bf F}\rangle_{\text{b}}$ and torque $\langle\Gamma\rangle_{\text{b}}$, where $\langle\cdot\rangle_{\text{b}}$ denotes an average at fixed object configuration, an instantaneous friction $\bm{\zeta}$, and a white noise $\bm{\xi}$. In this framework, the nine coefficients entering the friction matrix $\bm{\zeta}$ are given by an Agarwal formula Agarwal (1972)
@@
$\bm{\zeta}=\int_{0}^{\infty}\text{d}t\begin{bmatrix}\langle\delta\text{\bf F}(t)\bm{\nabla}_{\bm{R}}\ln\rho_{\text{b}}(0)\rangle_{\text{b}}&\langle\delta\text{\bf F}(t)\partial_{\Theta}\ln\rho_{\text{b}}(0)\rangle_{\text{b}}\\ \langle\delta\Gamma(t)\bm{\nabla}_{\bm{R}}^{T}\ln\rho_{\text{b}}(0)\rangle_{\text{b}}&\langle\delta\Gamma(t)\partial_{\Theta}\ln\rho_{\text{b}}(0)\rangle_{\text{b}}\end{bmatrix},$ (4)
@@
where $\delta\text{\bf F}=\text{\bf F}-\langle\text{\bf F}\rangle_{\text{b}}$ and $\delta\Gamma=\Gamma-\langle\Gamma\rangle_{\text{b}}$. The three Gaussian noises are characterized by their correlation matrix $\bm{\lambda}$, defined from
$\langle\bm{\xi}(t)\otimes\bm{\xi}(t^{\prime})\rangle=\bm{\lambda}\delta_{+}(t-t^{\prime})+\bm{\lambda}^{T}\delta_{-}(t-t^{\prime})\,,$ (5)
@@
where $\delta_{\pm}$ is the restriction of the Dirac function to $\mathbb{R}_{\pm}$ and $\bm{\lambda}$ satisfies the Green-Kubo formula
@@
$\bm{\lambda}=\int_{0}^{\infty}\text{d}t\begin{bmatrix}\langle\delta\text{\bf F}(t)\delta\text{\bf F}(0)\rangle_{\text{b}}&\langle\delta\text{\bf F}(t)\delta\Gamma(0)\rangle_{\text{b}}\\ \langle\delta\Gamma(t)\delta\text{\bf F}^{T}(0)\rangle_{\text{b}}&\langle\delta\Gamma(t)\delta\Gamma(0)\rangle_{\text{b}}\end{bmatrix}\,.$ (6)
@@
Figure 2: Confined disk: effective equilibrium and circulating momentum currents ($\ell_{p}=10$, $\ell_{g}=5$). (a) The disk’s potential (solid) and kinetic (dashed) energies converge to the Boltzmann distribution in the adiabatic limit ($M\in[10^{-2},10^{3}]$ increasing from light to dark). (b) Even and odd thermometers, constructed from $\bm{\lambda}$, $\bm{\zeta}$, $\mathbf{D}$, and $\bm{\mu}$, all converge to a single $T^{\text{eff}}$ in the adiabatic limit. (c) Probability density $\rho_{\bm{R}}$ (heatmap) and flux $\bm{J}_{\bm{R}}^{\text{ss}}$ (arrows) of the confined disk (cartoon overlay). Steady circulating currents (top) vanish at large $M$ (bottom) due to the odd Einstein relation (9). (d) $\rho_{\bm{P}}$ and $\bm{J}_{\bm{P}}^{\text{ss}}$ of the same system. Circulation persists even in the adiabatic limit. The direction of rotation, determined by competition between $\zeta_{\perp}$ and $\lambda_{\perp}$, reverses at smaller $M$.
@@
Equations (3)-(4) provide the most general dynamics for a rigid body in a chiral active bath, within the adiabatic approximation. They hold in any dimension $d\geq 2$ but neglect long-time tails due to conservation laws van Beijeren (1982); Dorfman et al. (2021), that can play an important role in low-dimensional active systems Granek et al. (2020). The object dynamics is then entirely defined by the sole knowledge of $\bm{\zeta}$, $\bm{\lambda}$, and the ratchet forces, $\langle\text{\bf F}\rangle_{\text{b}}$ and $\langle\Gamma\rangle_{\text{b}}$, which are determined by the object shape. We now apply these theoretical results to demonstrate how the object departs from an effective equilibrium limit as its symmetry decreases.
@@
The $SO(2)$ swirling disk. We begin by considering a circular disk, which has only translational degrees of freedom. By symmetry, $\langle\text{\bf F}\rangle_{\text{b}}={\bf 0}$, and the object dynamics reduce to
@@
The combined isotropy of the bath and the disk make $\bm{\zeta}$ and $\bm{\lambda}$ belong to the space of isotropic $2\times 2$ matrices, which is spanned by $\bm{\delta}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ and $\text{\bf A}=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}$, so that $\bm{\zeta}_{{\bm{P}}{\bm{P}}}=\zeta_{\parallel}\bm{\delta}+\zeta_{\perp}\text{\bf A}$ and $\bm{\lambda}_{{\bm{P}}{\bm{P}}}=\lambda_{\parallel}\bm{\delta}+\lambda_{\perp}\text{\bf A}$. Here, $\zeta_{\parallel}$ and $\lambda_{\parallel}$ are the usual friction and noise correlations, respectively. Their odd counterparts, $\zeta_{\perp}$ and $\lambda_{\perp}$, stem from the breaking of time-reversal and parity symmetry of the bath. As we show below, they allow for steady circulating currents that are the counterparts to those observed in confined chiral active matter van Zuiden et al. (2016); Souslov et al. (2017); Dasbiswas et al. (2018); Soni et al. (2019); Yang et al. (2021); Vuijk et al. (2020); Kalz et al. (2024); Vega Reyes et al. (2022); Caprini and Marconi (2019); Caprini et al. (2023); Li et al. (2024); Wittmann et al. (2024).
To investigate these odd currents, we confine the disk using a harmonic potential $U({\bm{R}})=\frac{1}{2}k|{\bm{R}}|^{2}$, as shown in Fig. 2. (Our results extend directly to more general potentials.) The simulations in Fig. 2a show how the nonequilibrium dynamics of light disks are replaced at large $M$ by an effective equilibrium one, leading to the Boltzmann steady-state solution of Eq. (7) given by
@@
remarkably holding for the odd parts as well as the even. Thus, while $D_{\perp}$ and $\mu_{\perp}$ can individually drive currents satisfying $\bm{\nabla}_{\bm{R}}\cdot\bm{J}_{\bm{R}}^{\text{ss}}=0$, these exactly cancel at large $M$, as shown in Fig. 2c (bottom panel), leading to an effective equilibrium apparently reminiscent to that observed for a bath of achiral particles Solon and Horowitz (2022). In contrast, when $M$ is small, the disk departs from the equilibrium distribution of Eq. (8). The competition between $D_{\perp}$ and $\mu_{\perp}$ then results in steady flows illustrated in Fig. 2c (top panel).
In the adiabatic limit, the flux of $\rho_{\bm{P}}$ is $\bm{J}^{\text{ss}}_{\bm{P}}=-\left(\bm{\lambda}-T^{\text{eff}}\bm{\zeta}\right)\bm{\nabla}_{\bm{P}}\rho_{\bm{P}}$. The continuity condition $\bm{\nabla}_{\bm{P}}\cdot\bm{J}^{\text{ss}}_{\bm{P}}=0$ requires that the even parts of $\bm{\lambda}$ and $\bm{\zeta}$ be related by a second fluctuation-dissipation theorem, but places no such requirement on the odd parts:
@@
The inequality in Eq. (10) results in odd currents that persist even in the adiabatic limit (Fig. 2d), and can even change direction as the disk mass is decreased. This is in sharp contrast with both chiral equilibrium baths, which require $\lambda_{\perp}=T^{\text{eff}}\zeta_{\perp}$, and achiral active ones Solon and Horowitz (2022).
We stress that Eqs (8)-(10) provide independent thermometers to measure the effective temperature of the system. As shown in Fig. 2b, they all agree in the adiabatic limit. All involved quantities remain well-defined even beyond this limit, and are of rheological interest Saintillan (2018); Shea et al. (2024), but cease to be directly connected through an effective temperature.
Figure 3: Effective temperatures of the spinning rod ($\ell_{\mathrm{rod}}=5$, $\ell_{p}=10$, and $\ell_{g}=5$). (a) The angular velocities of rods pinned at the center or free to translate agree at large $M$. (b) A rod confined to a harmonic potential exhibits rotational (blue) and translational (green) dynamics that independently recover equilibrium-like statistics at large $M$, with the mean potential (solid) and kinetic (dashed) energies agreeing, but at two distinct effective temperatures $T^{\text{eff}}_{\bm{R}}$ and $T^{\text{eff}}_{\Theta}$.
@@
The $C_{2}$ Spinning rod. We now proceed beyond the isotropic disk by considering a passive rod, pictured in Fig. 1b. While reflection symmetry is lost due to the bath chirality, the system maintains a two-fold rotational symmetry ($C_{2}$). Consequently, $\langle\text{\bf F}\rangle_{\text{b}}=0$, a result that holds for any object with $n$-fold rotational symmetry ($C_{n}$). The bath chirality nevertheless drives asymmetric accumulation around the rod, leading to a net torque and a rotational ratchet behavior.
Furthermore, Eqs. (6)-(4) show the elements of $\bm{\zeta}$ and $\bm{\lambda}$ that couple rotational and translational motion to vanish under $C_{n}$ symmetry. This decoupling means that the translational dynamics of the rod are structurally identical to those of the isotropic disk, while the rotational dynamics obeys
@@
Another consequence of the decoupling is that translational and rotational degrees of freedom are both equilibrated, but at different temperatures. The stationary distribution of a rod in a confining potential $U=\frac{1}{2}k|{\bm{R}}|^{2}+\frac{1}{2}k_{\Theta}\Theta^{2}$ is
@@
where $\delta\Theta=\Theta-\langle\Gamma\rangle_{\text{b}}/k_{\Theta}$. As shown in Fig. 3b, the translational temperature $T^{\text{eff}}_{\bm{R}}$ is nearly twice as “hot” as $T^{\text{eff}}_{\Theta}$, even in the adiabatic limit, where the translational and rotational degrees of freedom individually obey equipartition between kinetic and potential energies.
We note that the odd circulation in momentum space of the disk is associated with a nonzero entropy production rate, which diverges as $\sigma\propto 1/\epsilon$. Similarly, the ratchet rotation of the rod leads to $\sigma=\frac{\langle\Gamma\rangle_{\text{b}}\Omega}{T^{\text{eff}}_{\Theta}}$. These results seem to contradict the equilibrium-like statistics of Eq. (8) for the disk and of Eq. (11) for the rod. However, this paradox is resolved by noting that statistical reversibility is restored for objects with $C_{n}$ symmetries under combined time reversal and chirality inversion ($\omega_{0}\rightarrow-\omega_{0}$). As we now show, this is entirely lost when the object symmetry is further reduced.

### v2

@@ Abstract
When submerged in a chiral active bath, a passive object becomes a spinning ratchet imbued with odd transport properties. We present the most general Langevin dynamics for a rigid body in a chiral active bath, in the adiabatic limit of large object mass. For rotationally symmetric objects, odd diffusion and odd mobility are connected by an Einstein relation, that we show numerically to break down outside the adiabatic limit. As the object symmetry decreases, its dynamics becomes increasingly irreversible: a massive disk exhibits an effective equilibrium dynamics, while a rod admits distinct translational and rotational temperatures, and a wedge is fully irreversible. Conversely, this departure from equilibrium can be read in universal far-field currents and density modulations of the bath, which we measure numerically and derive analytically.
@@
A microscopic object in a fluid bath inherits its dynamics from collisions with the bath particles. When the bath is in equilibrium, the object obeys the Einstein relation and exhibits Boltzmann statistics. Active baths, by contrast, allow a richer set of phenomena and have thus attracted a lot of attention Wu and Libchaber (2000); Loi et al. (2008); Underhill et al. (2008); Leptos et al. (2009); Dunkel et al. (2010); Kurtuldu et al. (2011); Foffano et al. (2012); Miño et al. (2013); Morozov and Marenduzzo (2014); Burkholder and Brady (2017); Chaki and Chakrabarti (2018); Kanazawa et al. (2020); Knežević and Stark (2020); Reichert et al. (2021); Maes (2020); Granek et al. (2022); Solon and Horowitz (2022); Sorkin et al. (2024). Notable phenomena include ratchet motions Magnasco (1993); Van den Broeck et al. (2004); Angelani et al. (2011); Gnoli et al. (2013); Reichhardt and Reichhardt (2017), demonstrating how an active bath can power microscopic engines Krishnamurthy et al. (2016); Zakine et al. (2017); Pietzonka et al. (2019); Fodor and Cates (2021). When asymmetric gears are inserted into an active bacterial bath Di Leonardo et al. (2010); Sokolov et al. (2010); Anand et al. (2024), rotational ratchet motion arises due to broken time-reversal symmetry of the bath and broken parity symmetry (i.e. chirality) of the object.
@@
More recent investigations have looked at chiral active baths, in which both of these symmetries are broken by the bath itself. In such baths, even symmetric objects become rotational ratchets Li et al. (2023) imbued with odd diffusivity Hargus et al. (2021); Kalz et al. (2022) and odd mobility Reichhardt and Reichhardt (2019); Poggioli and Limmer (2023). While extensive analytical and numerical results are available for achiral baths, little is known theoretically of the dynamics of passive objects embedded in chiral active fluids. The latter are however abundant in both biological Drescher et al. (2009); Petroff et al. (2015); Beppu et al. (2021); Tan et al. (2022) and synthetic active systems Soni et al. (2019); Bililign et al. (2022). This raises two key questions: how are the emerging properties of the object (and the bath) interrelated, and how are they influenced by the object shape?
@@
In this Letter, we answer these questions by constructing a general theory to describe the dynamics of a massive object embedded in a chiral active bath. This allows us to show that the object dynamics become increasingly rich as its symmetry decreases. Specifically, we consider the disk, rod, and wedge depicted in Fig. 1. For the disk, we find an effective equilibrium description for the translational degrees of freedom, accompanied by circulating currents in momentum space. The rod departs from this effective equilibrium, by displaying distinct temperatures for the rotational and translational degrees of freedom. Last, the wedge is fully irreversible, behaving as both a translational and rotational ratchet. Finally, we show how the object symmetry influences the bath itself through the generation of long-range density modulation and currents. Here we focus on the presentation of our results; numerical details are provided in the End Matter while detailed derivations can be found in a companion paper Hargus et al. .
Figure 1: Passive objects inherit chiral dynamics from the bath. A disk (a) exhibits odd diffusivity $D_{\perp}$, a rod (b) additionally rotates with a ratchet velocity $\Omega$, and a passive wedge (c) additionally translates with ratchet velocities $v_{\parallel}$ and $v_{\perp}$.
@@
$M\ddot{\bm{R}}=\dot{\bm{P}}={\bm{F}}\,,\quad I\ddot{\Theta}=\dot{L}=\Gamma\,,$ (1)
@@
where ${\bm{F}}$ and $\Gamma$ are the total force and torque due to the bath particles. The latter are modeled as chiral active Brownian particles, whose positions $\bm{r}_{i}$ and orientations $\theta_{i}$ evolve as
$\gamma\dot{\bm{r}}_{i}={\bm{F}}_{i}+f_{0}{\bm{u}}_{i}\,,\quad\dot{\theta}_{i}=\omega_{0}+\sqrt{2D_{r}}\xi_{i}\,.$ (2)
@@
Here, $\gamma$ is the substrate friction, $f_{0}$ is the active force oriented along ${\bm{u}}_{i}=[\cos(\theta_{i}),\sin(\theta_{i})]^{T}$, $\omega_{0}$ is an angular drift, $D_{r}$ is the rotational diffusivity, and $\xi_{i}$ are centered unit-variance Gaussian white noises. The bath displays two active lengthscales: the persistence length $\ell_{p}=\frac{f_{0}}{D_{r}\gamma}$ and gyroradius $\ell_{g}=\frac{f_{0}}{|\omega_{0}|\gamma}$. Object-bath interactions are reciprocal, with $\sum_{i}{\bm{F}}_{i}=-{\bm{F}}$.
@@
If the motion of the object is very slow relative to that of the bath, we can work in the adiabatic limit Van Kampen and Oppenheim (1986); Granek et al. (2020); Solon and Horowitz (2022); Jayaram and Speck (2023) where the probability distribution of the full system factorizes as $\rho({\bm{R}},{\bm{P}},\Theta,L,{\bm{r}}^{N},{\bm{u}}^{N},t)\approx\rho_{\text{o}}({\bm{R}},{\bm{P}},\Theta,L,t)\rho_{\text{b}}({\bm{r}}^{N},{\bm{u}}^{N}|{\bm{R}},\Theta)$. This leaves all time dependence in $\rho_{\text{o}}$, while the bath relaxes to the instantaneous object configuration. This separation of timescales is characterized by the dimensionless parameter $\epsilon=\sqrt{\gamma/D_{r}M}$, which is small in the limit of massive objects. Following established procedures Van Kampen and Oppenheim (1986); Solon and Horowitz (2022), we integrate out the bath degrees of freedom to obtain underdamped Langevin dynamics for the object:
@@
$\begin{split}\begin{bmatrix}\dot{\bm{P}}\\ \dot{L}\end{bmatrix}&=\begin{bmatrix}\langle{\bm{F}}\rangle_{\text{b}}\\ \langle\Gamma\rangle_{\text{b}}\end{bmatrix}-\underbrace{\begin{bmatrix}\bm{\zeta}_{{\bm{P}}{\bm{P}}}&\bm{\zeta}_{{\bm{P}}L}\\ \bm{\zeta}_{L{\bm{P}}}&\zeta_{LL}\end{bmatrix}}_{\bm{\zeta}}\begin{bmatrix}\frac{1}{M}{\bm{P}}\\ \frac{1}{I}L\end{bmatrix}+\underbrace{\begin{bmatrix}\bm{\xi}_{P}(t)\\ \xi_{L}(t)\end{bmatrix}}_{\bm{\xi}}\,.\end{split}$ (3)
@@
The bath produces three effects on the object: a mean force $\langle{\bm{F}}\rangle_{\text{b}}$ and torque $\langle\Gamma\rangle_{\text{b}}$ (where $\langle\cdot\rangle_{\text{b}}$ denotes an average at fixed object configuration), an instantaneous friction $\bm{\zeta}$, and Gaussian noise $\bm{\xi}$. In this framework, the nine coefficients entering the friction matrix $\bm{\zeta}$ are given by an Agarwal formula Agarwal (1972)
@@
$\bm{\zeta}=\int_{0}^{\infty}\text{d}t\begin{bmatrix}\langle\delta{\bm{F}}(t)\bm{\nabla}_{\bm{R}}\ln\rho_{\text{b}}(0)\rangle_{\text{b}}&\langle\delta{\bm{F}}(t)\partial_{\Theta}\ln\rho_{\text{b}}(0)\rangle_{\text{b}}\\ \langle\delta\Gamma(t)\bm{\nabla}_{\bm{R}}^{T}\ln\rho_{\text{b}}(0)\rangle_{\text{b}}&\langle\delta\Gamma(t)\partial_{\Theta}\ln\rho_{\text{b}}(0)\rangle_{\text{b}}\end{bmatrix},$ (4)
@@
where $\delta{\bm{F}}={\bm{F}}-\langle{\bm{F}}\rangle_{\text{b}}$ and $\delta\Gamma=\Gamma-\langle\Gamma\rangle_{\text{b}}$. The Gaussian noises are characterized by their correlation matrix $\bm{\lambda}$, defined by
$\langle\bm{\xi}(t)\bm{\xi}(t^{\prime})\rangle=\bm{\lambda}\delta_{+}(t-t^{\prime})+\bm{\lambda}^{T}\delta_{-}(t-t^{\prime})\,,$ (5)
@@
where $\delta_{\pm}$ is the restriction of the Dirac function to $\mathbb{R}_{\pm}$ Chun et al. (2018) and $\bm{\lambda}$ satisfies the Green-Kubo formula
@@
$\bm{\lambda}=\int_{0}^{\infty}\text{d}t\begin{bmatrix}\langle\delta{\bm{F}}(t)\delta{\bm{F}}(0)\rangle_{\text{b}}&\langle\delta{\bm{F}}(t)\delta\Gamma(0)\rangle_{\text{b}}\\ \langle\delta\Gamma(t)\delta{\bm{F}}^{T}(0)\rangle_{\text{b}}&\langle\delta\Gamma(t)\delta\Gamma(0)\rangle_{\text{b}}\end{bmatrix}\,.$ (6)
@@
Figure 2: Confined disk: effective equilibrium and circulating currents ($\ell_{p}=10$, $\ell_{g}=5$). (a) The disk’s potential (solid) and kinetic (dashed) energies converge to the Boltzmann distribution in the adiabatic limit ($M\in[10^{-2},10^{3}]$ increasing from light to dark). (b) Even and odd thermometers, constructed from $\bm{\lambda}$, $\bm{\zeta}$, $\mathbf{D}$, and $\bm{\mu}$, all converge to a single $T^{\text{eff}}$ in the adiabatic limit. (c) Probability density $\rho_{\bm{R}}$ (heatmap) and flux $\bm{J}_{\bm{R}}^{\text{ss}}$ (arrows) of the confined disk (cartoon overlay). Steady circulating currents (top) vanish at large $M$ (bottom) due to the odd Einstein relation (9). (d) $\rho_{\bm{P}}$ and $\bm{J}_{\bm{P}}^{\text{ss}}$ of the same system. Circulation persists even in the adiabatic limit. The direction of rotation, determined by competition between $\zeta_{\perp}$ and $\lambda_{\perp}$, reverses at smaller $M$.
@@
Equations (3)-(6) provide the most general dynamics for a rigid body in a chiral active bath, within the adiabatic approximation. They hold in any dimension $d\geq 2$ but neglect long-time tails due to conservation laws van Beijeren (1982); Dorfman et al. (2021), that can play an important role in low-dimensional active systems Granek et al. (2020). The object dynamics is then entirely determined by $\bm{\zeta}$, $\bm{\lambda}$, $\langle{\bm{F}}\rangle_{\text{b}}$ and $\langle\Gamma\rangle_{\text{b}}$, which are set by the object shape. We now apply these theoretical results to show how the object departs from an effective equilibrium limit as its symmetry decreases.
@@
The $SO(2)$ swirling disk. We begin by considering a circular disk, which has only translational degrees of freedom. By symmetry, $\langle{\bm{F}}\rangle_{\text{b}}={\bf 0}$, and Eq. (3) reduces to
@@
The combined isotropy of the bath and the disk make $\bm{\zeta}$ and $\bm{\lambda}$ belong to the space of isotropic $2\times 2$ matrices, which is spanned by $\mathbf{1}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ and $\text{\bf A}=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}$, so that $\bm{\zeta}_{{\bm{P}}{\bm{P}}}=\zeta_{\parallel}\mathbf{1}+\zeta_{\perp}\text{\bf A}$ and $\bm{\lambda}_{{\bm{P}}{\bm{P}}}=\lambda_{\parallel}\mathbf{1}+\lambda_{\perp}\text{\bf A}$. Here, $\zeta_{\parallel}$ and $\lambda_{\parallel}$ are the usual friction and noise correlations, respectively. Their odd counterparts, $\zeta_{\perp}$ and $\lambda_{\perp}$, stem from the breaking of time-reversal and parity symmetry of the bath Hargus et al. (2025). As we show below, these drive steady circulating currents of the object, providing a counterpoint to those observed in confined chiral active fluids van Zuiden et al. (2016); Souslov et al. (2017); Dasbiswas et al. (2018); Soni et al. (2019); Yang et al. (2021); Vuijk et al. (2020); Kalz et al. (2024); Vega Reyes et al. (2022); Caprini and Marconi (2019); Caprini et al. (2023); Li et al. (2024); Wittmann et al. (2024).
To investigate these odd currents, we confine the disk using a harmonic potential $U({\bm{R}})=\frac{1}{2}k|{\bm{R}}|^{2}$, as shown in Fig. 2. The simulations in Fig. 2a show how the nonequilibrium dynamics of light disks are replaced at large $M$ by an effective equilibrium one, leading to the Boltzmann steady-state solution of Eq. (7) given by
@@
remarkably holding for the odd parts as well as the even. Thus, while $D_{\perp}$ and $\mu_{\perp}$ can individually drive currents satisfying $\bm{\nabla}_{\bm{R}}\cdot\bm{J}_{\bm{R}}^{\text{ss}}=0$, these exactly cancel at large $M$, as shown in Fig. 2c (bottom panel), leading to an effective equilibrium reminiscent of that observed for a bath of achiral active particles Solon and Horowitz (2022). In contrast, when $M$ is small, the disk departs from the equilibrium distribution of Eq. (8). Competition between $D_{\perp}$ and $\mu_{\perp}$ then results in steady flows illustrated in Fig. 2c (top panel).
Figure 3: Ratchet motion and effective temperatures of the spinning rod ($\ell_{\mathrm{rod}}=5$, $\ell_{p}=10$, and $\ell_{g}=5$). (a) The angular velocities of rods pinned at the center or free to translate agree at large $M$. (b) A rod confined to a harmonic potential exhibits rotational (blue) and translational (green) dynamics that independently recover equilibrium-like statistics at large $M$, with the mean potential (solid) and kinetic (dashed) energies agreeing, but at two distinct effective temperatures $T^{\text{eff}}_{\bm{R}}$ and $T^{\text{eff}}_{\Theta}$.
In the adiabatic limit, the flux of $\rho_{\bm{P}}$ is $\bm{J}^{\text{ss}}_{\bm{P}}=\left(T^{\text{eff}}\bm{\zeta}-\bm{\lambda}\right)\bm{\nabla}_{\bm{P}}\rho_{\bm{P}}$. The continuity condition $\bm{\nabla}_{\bm{P}}\cdot\bm{J}^{\text{ss}}_{\bm{P}}=0$ requires that the even parts of $\bm{\lambda}$ and $\bm{\zeta}$ be related by a second fluctuation-dissipation theorem, but places no such requirement on the odd parts:
@@
The inequality in Eq. (10) results in odd currents that persist even in the adiabatic limit (Fig. 2d). This mismatch is a unique odd signature of chiral nonequilibrium baths.
We stress that Eqs (8)-(10) provide independent thermometers to measure the effective temperature of the system (Fig. 2b). Beyond the adiabatic regime, all involved quantities remain well-defined and are of rheological interest Saintillan (2018); Shea et al. (2024), but cease to be directly connected through an effective temperature.
@@
The $C_{2}$ spinning rod. We now proceed beyond the isotropic disk to consider the rod pictured in Fig. 1b. The chiral bath now drives asymmetric accumulation around the object, leading to a net torque. The full system, however, maintains two-fold rotational symmetry ($C_{2}$). Consequently, $\langle{\bm{F}}\rangle_{\text{b}}=0$, a result that holds for any object with $n$-fold rotational symmetry ($C_{n}$).
Furthermore, Eqs. (4) and (6) show the elements of $\bm{\zeta}$ and $\bm{\lambda}$ that couple rotational and translational motion to vanish under $C_{n}$ symmetry. This decoupling means that the translational dynamics of the rod are structurally identical to those of the isotropic disk, while the rotational dynamics obeys
@@
Another consequence of this decoupling is that translational and rotational degrees of freedom are independently equilibrated, but at different temperatures. The stationary distribution of a rod in a confining potential $U=\frac{1}{2}k|{\bm{R}}|^{2}+\frac{1}{2}k_{\Theta}\Theta^{2}$ is
@@
where $\delta\Theta=\Theta-\langle\Gamma\rangle_{\text{b}}/k_{\Theta}$. As shown in Fig. 3b, the translational temperature $T^{\text{eff}}_{\bm{R}}$ is nearly twice as “hot” as $T^{\text{eff}}_{\Theta}$.
The effective equilibrium observed for both the disk and rod can be rationalized by noting that the adiabatic dynamics of rotationally-symmetric objects are statistically reversible under combined time reversal and chirality inversion ($\omega_{0}\rightarrow-\omega_{0}$argus et al. . This hidden symmetry is lost when $C_{n}$ symmetry is broken, as for the wedge studied in the next section.